Find the length of the curve given by x = t
1
Expert's answer
2021-09-30T08:49:33-0400

length of the curve,

"L=\int_{t_1}^{t_2}\sqrt{(\frac{dx}{dt})^2+(\frac{dy}{dt})^2}dt"

by using above formula, we get

"L=\int_{t_1}^{t_2}\sqrt{(1)^2+(0)^2}dt\\
=\int_{t_1}^{t_2}dt\\
=t_2-t_1"